# A Survey of Vibe Coding with Large Language Models

### v1

@@ 9.2 Code Reliability and Security in Vibe Coding
The core trade-off of Vibe Coding is speed versus certainty. By abstracting away line-by-line code authoring, it introduces a fundamental challenge to ensuring code reliability and security. When a developer validates code based on its "vibe" or observed outcome, they may inadvertently approve implementations containing subtle but critical flaws, from resource leaks and race conditions to severe security vulnerabilities. This risk is amplified because LLMs are trained on vast corpora of public code, which inevitably includes buggy and insecure examples [584, 585, 586]. The model may therefore reproduce these vulnerabilities in novel contexts. Consequently, for Vibe Coding to be a viable methodology for production-grade software, it must be augmented by a new generation of automated guardrails that operate continuously and intelligently within the development loop.
@@ 9.2.1 The Inadequacy of Manual Review
Traditional security assurance relies heavily on manual expert code review [587]. This practice is fundamentally incompatible with the philosophy of Vibe Coding. It is paradoxical to gain development speed by having an AI generate code, only to lose that speed by requiring a human to manually inspect every line of the output. Furthermore, the developer prompting the AI may lack the specific security expertise to spot vulnerabilities in the generated code, especially if it uses an unfamiliar library or a complex algorithm. The very premise of Vibe Coding—trusting the agent to handle implementation details—necessitates a shift from human-centric inspection to automated, real-time validation.
@@ 9.2.2 Architecting an Integrated Security and Reliability Feedback Loop
The feedback loop should begin even before code is generated. An intelligent agent, aware of the project’s security requirements, can analyze the developer’s prompt for security-sensitive keywords (e.g., "authentication," "file upload," "SQL query"). This can trigger the agent to inject a security-specific context, such as secure coding guidelines or templates, to steer the LLM towards generating more robust code from the outset [588].
@@
As the LLM streams its generated code into the developer’s IDE, a real-time SAST engine should analyze it on the fly. This is a significant evolution from traditional SAST tools, which often require complete, compilable code [589]. New, AI-enhanced SAST tools are emerging that can parse incomplete code snippets and use machine learning to identify potential vulnerabilities with fewer false positives [590]. Platforms like Amazon CodeWhisperer, with its integrated security scan feature [591] represent an early step in this direction. A mature system would not just flag an issue but could feed the vulnerability information back to the LLM in a closed loop, prompting it to self-correct its own output before the developer even sees the flawed version.
@@
The "validate" step of the Vibe Coding loop is a critical intercept point. When the developer executes the generated code to check its "vibe," the execution should occur within an instrumented sandbox. This environment can perform just-in-time DAST, monitoring for runtime errors like memory leaks, uncaught exceptions, and insecure network communications. It could also perform automated fuzzing on newly generated API endpoints or data processing functions to test for edge-case vulnerabilities, providing immediate feedback on the code’s dynamic behavior. The integration of static and dynamic analysis offers a complementary approach to cover a wider range of potential issues [592, 593].
@@ 9.2.3 The Human-in-the-Loop as a Final Arbiter
Even with this sophisticated automation, the human developer remains the crucial arbiter. The role of the developer is not eliminated but elevated. They are responsible for interpreting the outputs of these automated security tools, resolving ambiguities, and making the final risk-based decision. The goal of the tooling is to empower the developer with immediate, actionable intelligence, transforming the security process from a separate, delayed phase into an intrinsic property of the code generation act itself. However, the technical specifications and architectural blueprints for these deeply integrated, real-time security frameworks in Vibe Coding workflows are not yet standardized [590, 20] and building them remains a major challenge for the research and tool-development communities. Without solving this, Vibe Coding risks being confined to prototyping and non-critical applications, unable to deliver on its promise for reliable, secure, and production-ready software systems.
@@ 9.3 Scalable Oversight of Vibe Coding Agents
As coding agents evolve toward autonomous code generation and deployment [243, 244], the scope of oversight must expand from localized code verification to system-level governance. While Section 9.2 examined reliability and security at the level of generated code, the next frontier concerns the supervision of agents that operate continuously across repositories, pipelines, and environments. The central challenge lies in scaling assurance mechanisms to match the increasing autonomy, speed, and operational complexity of these agents. Traditional safeguards designed for human-authored code, for example, peer review [572, 594, 595], and post-deployment auditing [596, 597], cannot feasibly monitor the millions of tokens or decisions that modern agents can generate per hour. This growing disparity between agent capability and human supervision capacity has become a critical obstacle to safe and accountable Vibe Coding workflows.
@@ 9.3.1 Emerging Risks of Vide Coding Workflow
Recent empirical analyses show that the introduction of autonomous agents into production pipelines correlates with a tenfold increase in security warnings and technical debt accumulation within six months of adoption [184, 584]. These findings reveal that human-centric verification processes fail to scale in environments where agents autonomously generate, refactor, and redeploy code. The challenge is not simply one of quality assurance but of maintaining interpretability and control in increasingly opaque decision pathways. Over-reliance on autonomous coding agents introduces new classes of risks:
@@
Cascading errors occur when autonomous agents propagate flawed outputs through interconnected software pipelines, transforming local inaccuracies into systemic failures. In multi-agent coding environments, a single agent’s erroneous completion, such as generating insecure logic or malformed interfaces, can trigger downstream faults as subsequent agents consume and redeploy its outputs. [598] demonstrate that even minor perturbations from “faulty” agents can destabilize collaborative workflows, revealing the fragility of LLM-based coordination without strong fault isolation. Similarly, [122] observe that emergent interactions in AgentMesh amplify defects across code generation stages, while [599] characterize “butterfly effects” in LLM toolchains, where trivial parameter-filling mistakes escalate into broad execution failures. Extending this view, [600] find that scientific discovery agents exhibit correlated errors due to shared model assumptions and data dependencies. Collectively, these studies underscore that LLM-mediated ecosystems lack the containment and traceability necessary to prevent error propagation—rendering traditional debugging and peer review inadequate in fully autonomous settings.
@@
Dependency proliferation denotes the uncontrolled expansion of software and library linkages arising from agent-driven code generation and refactoring. Autonomous coding agents frequently introduce or hallucinate dependencies beyond human oversight, compounding security and maintainability challenges. [601] reveal that nearly one-fifth of packages suggested by code-generation models are nonexistent or untrusted, creating opportunities for dependency-confusion attacks. [602] identify similar propagation patterns in automated import suggestions, where vulnerable or deprecated modules reappear across independent tasks. Furthermore, [603] show that fuzz-testing agents inadvertently inject unverified binaries during test synthesis, and [604] demonstrate that build-automation agents routinely bypass provenance enforcement to optimize throughput. These behaviors collectively inflate the software’s attack surface and contravene emerging SBOM (Software Bill of Materials) and supply-chain transparency requirements. Dependency proliferation thus represents a structural form of technical debt—an accreting web of opaque linkages that undermines the long-term security and interpretability of autonomous development environments.
@@
Alignment failures emerge when an autonomous coding agent’s behavior diverges from the developer’s intent or the organization’s governance norms [605]. Unlike deterministic compilers, agentic systems interpret underspecified objectives probabilistically, optimizing for textual success rather than normative correctness. [606] empirically show that goal-oriented coding agents often maximize explicit task metrics while disregarding implicit safety or maintainability constraints embedded in natural-language prompts. Complementarily, [607] describe “agentic misalignment” as an emergent property of reinforcement-trained code agents that overfit reward structures, leading to functionally correct yet normatively undesirable behaviors. Such divergence erodes accountability: once decisions are embedded in opaque model reasoning, tracing responsibility between human and agent becomes ambiguous. Alignment failures therefore epitomize the epistemic opacity of autonomous software systems, emphasizing the necessity of interpretable reward modeling, continuous supervision, and human-in-the-loop intervention to ensure fidelity between developer intent and agent execution.
Addressing these systemic risks therefore necessitates a layered oversight architecture that integrates automated analysis, formal verification, and human-in-the-loop governance. Modern static and dynamic analysis pipelines now incorporate model-aware scanners capable of parsing incomplete code fragments and predicting vulnerabilities in real time [509, 501], extending beyond conventional SAST [608] and DAST [609] frameworks by interfacing directly with generation streams so that feedback can be injected before completion. Complementing these, formal verification frameworks such as Kani [610] and ACCA [611] demonstrate that lightweight proofs of functional correctness can be applied to AI-generated modules, though scalability remains limited. Reinforcement-learning-based watchdog agents offer a parallel safeguard: autonomous monitors trained to detect deviations from project constraints, perform behavioral tracing, and intervene when agents exceed operational bounds [612, 613]. Collectively, these technologies lay the groundwork for continuous, adaptive oversight that aligns autonomous coding processes with human accountability and system safety.
@@ 9.3.2 Toward Scalable Oversight Architectures
Achieving effective oversight in autonomous coding environments requires architectures that scale alongside agent capabilities. A central idea in recent alignment research is enabling weak-to-strong generalization, using relatively weak supervisors (humans or smaller models) to robustly guide much more powerful coding agents [614]. In contrast to traditional settings where humans directly review weaker AI outputs, future agents may produce “millions of lines of novel and potentially dangerous code” beyond any one human’s full comprehension. Scalable oversight architectures address this gap by bootstrapping supervision: leveraging layered AI assistants and automated checks so that even limited feedback can generalize into reliable constraints on agent behavior [615]. Alignment strategies such as iterated amplification, recursive reward modeling, and AI debate all share this goal, allowing networks of weaker reviewers, often AIs themselves, to collectively evaluate and correct stronger agents’ decisions [616, 617]. This paradigm of “weak-to-strong oversight” underpins emerging proposals for governing advanced code-generation systems, ensuring that each increase in agent autonomy is met with a proportional increase in oversight capacity.
@@
One approach is to organize oversight in tiers, where an automated or human overseer supervises the coding agent at a high level, while that overseer is itself aided by subordinate AI tools. This nested oversight or recursive supervision strategy allows complex judgments to be decomposed into simpler subtasks manageable by weaker supervisors [615]. Recent studies demonstrate that a weak teacher model’s guidance can be amplified through ensemble feedback and iterative training such that a much stronger student model aligns with the teacher’s underlying intent [618]. OpenAI’s superalignment experiments showed that fine-tuning GPT-4 under the noisy guidance of a GPT-2-level model, augmented with confidence-based loss functions, could recover much of GPT-4’s original performance, achieving results comparable to GPT-3.5 on complex reasoning tasks [614]. These findings suggest that, with appropriate training and oversight scaffolds, coding agents can generalize from weak supervision to strong, aligned behavior. However, theory also indicates that scalability has limits: oversight success rates decline sharply when an agent’s capability surpasses its overseer by several hundred Elo points [615], underscoring the need for continuous co-evolution between overseers and agents.
@@
Another branch of scalable oversight employs multi-agent debate frameworks [434, 619, 620], where coding agents critique each other’s code contributions to expose errors or unsafe behaviors [621, 617]. Even a weaker human or AI judge can often identify the more truthful debater if the setup incentivizes evidence-based argumentation [617]. In the software domain, DebateCoder [622] applies this principle to code generation: two large language models alternately produce and test code while generating adversarial unit tests against each other. Each iteration forces self-correction until both code and tests converge to correctness. This mechanism fuses automated testing and AI critique, forming a self-correcting oversight loop that improves code quality and reliability without human intervention. Such multi-agent oversight reduces dependence on a single reviewer and increases the likelihood that at least one agent detects subtle flaws or misalignments in autonomously written code.
@@
Scalable oversight also requires continuous, model-aware monitoring embedded throughout the software pipeline. Model-integrated static analyzers and runtime scanners can evaluate partially generated code, flagging vulnerabilities or policy violations in real time [509, 501]. Reinforcement learning–based watchdog agents extend this capacity: autonomous monitors trained to detect deviations from specification and intervene or alert humans when the coding agent exceeds its operational bounds [612, 613].
Scalable oversight architectures for Vibe Coding agents converge on the principle of amplifying limited oversight into broad governance. Through hierarchical weak-to-strong supervision, multi-agent critique, and automated watchdogs, researchers aim to generalize weak feedback signals into strong preventive constraints that evolve with agent capability [618, 614, 615]. The challenge lies not only in constructing these systems but in maintaining their reliability as agents advance beyond direct human comprehension. By incrementally upgrading oversight tools alongside the agents themselves, the field moves toward an adaptive architecture, one where even superhuman coding systems remain answerable to human-aligned constraints.

### v2

@@ 9.2 Code Reliability and Security in Vibe Coding
The core trade-off of Vibe Coding is speed versus certainty. By abstracting away line-by-line code authoring, it introduces a fundamental challenge to ensuring code reliability and security. When a developer validates code based on its "vibe" or observed outcome, they may inadvertently approve implementations containing subtle but critical flaws, from resource leaks and race conditions to severe security vulnerabilities. This risk is amplified because LLMs are trained on vast corpora of public code, which inevitably includes buggy and insecure examples [pearce2025asleep, fu2023security, Perry2022AIInsecureCode]. The model may therefore reproduce these vulnerabilities in novel contexts. Consequently, for Vibe Coding to be a viable methodology for production-grade software, it must be augmented by a new generation of automated guardrails that operate continuously and intelligently within the development loop.
@@ 9.2.1 The Inadequacy of Manual Review
Traditional security assurance relies heavily on manual expert code review [Hannah2020RiskAssessment]. This practice is fundamentally incompatible with the philosophy of Vibe Coding. It is paradoxical to gain development speed by having an AI generate code, only to lose that speed by requiring a human to manually inspect every line of the output. Furthermore, the developer prompting the AI may lack the specific security expertise to spot vulnerabilities in the generated code, especially if it uses an unfamiliar library or a complex algorithm. The very premise of Vibe Coding—trusting the agent to handle implementation details—necessitates a shift from human-centric inspection to automated, real-time validation.
@@ 9.2.2 Architecting an Integrated Security and Reliability Feedback Loop
The feedback loop should begin even before code is generated. An intelligent agent, aware of the project’s security requirements, can analyze the developer’s prompt for security-sensitive keywords (e.g., "authentication," "file upload," "SQL query"). This can trigger the agent to inject a security-specific context, such as secure coding guidelines or templates, to steer the LLM towards generating more robust code from the outset [Res2024CopilotPromptSecurity].
@@
As the LLM streams its generated code into the developer’s IDE, a real-time SAST engine should analyze it on the fly. This is a significant evolution from traditional SAST tools, which often require complete, compilable code [Medeiros2020VulnerableDetection]. New, AI-enhanced SAST tools are emerging that can parse incomplete code snippets and use machine learning to identify potential vulnerabilities with fewer false positives [Chittibala2024AutomatedScanning]. Platforms like Amazon CodeWhisperer, with its integrated security scan feature [Pudari2023CopilotToPilot] represent an early step in this direction. A mature system would not just flag an issue but could feed the vulnerability information back to the LLM in a closed loop, prompting it to self-correct its own output before the developer even sees the flawed version.
@@
The "validate" step of the Vibe Coding loop is a critical intercept point. When the developer executes the generated code to check its "vibe," the execution should occur within an instrumented sandbox. This environment can perform just-in-time DAST, monitoring for runtime errors like memory leaks, uncaught exceptions, and insecure network communications. It could also perform automated fuzzing on newly generated API endpoints or data processing functions to test for edge-case vulnerabilities, providing immediate feedback on the code’s dynamic behavior. The integration of static and dynamic analysis offers a complementary approach to cover a wider range of potential issues [Aggarwal2006StaticDynamic, Keromytis2011MINESTRONE].
@@ 9.2.3 The Human-in-the-Loop as a Final Arbiter
Even with this sophisticated automation, the human developer remains the crucial arbiter. The role of the developer is not eliminated but elevated. They are responsible for interpreting the outputs of these automated security tools, resolving ambiguities, and making the final risk-based decision. The goal of the tooling is to empower the developer with immediate, actionable intelligence, transforming the security process from a separate, delayed phase into an intrinsic property of the code generation act itself. However, the technical specifications and architectural blueprints for these deeply integrated, real-time security frameworks in Vibe Coding workflows are not yet standardized [Chittibala2024AutomatedScanning, sapkota2025vibe] and building them remains a major challenge for the research and tool-development communities. Without solving this, Vibe Coding risks being confined to prototyping and non-critical applications, unable to deliver on its promise for reliable, secure, and production-ready software systems.
@@ 9.3 Scalable Oversight of Vibe Coding Agents
As coding agents evolve toward autonomous code generation and deployment [dong2025survey, wang2025ai], the scope of oversight must expand from localized code verification to system-level governance. While Section 9.2 examined reliability and security at the level of generated code, the next frontier concerns the supervision of agents that operate continuously across repositories, pipelines, and environments. The central challenge lies in scaling assurance mechanisms to match the increasing autonomy, speed, and operational complexity of these agents. Traditional safeguards designed for human-authored code, for example, peer review [rigby2013convergent, bacchelli2013expectations, mcintosh2014impact], and post-deployment auditing [dissanayake2022software, lin2007autopag], cannot feasibly monitor the millions of tokens or decisions that modern agents can generate per hour. This growing disparity between agent capability and human supervision capacity has become a critical obstacle to safe and accountable Vibe Coding workflows.
@@ 9.3.1 Emerging Risks of Vide Coding Workflow
Recent empirical analyses show that the introduction of autonomous agents into production pipelines correlates with a tenfold increase in security warnings and technical debt accumulation within six months of adoption [barke2023grounded, pearce2025asleep]. These findings reveal that human-centric verification processes fail to scale in environments where agents autonomously generate, refactor, and redeploy code. The challenge is not simply one of quality assurance but of maintaining interpretability and control in increasingly opaque decision pathways. Over-reliance on autonomous coding agents introduces new classes of risks:
@@
Cascading errors occur when autonomous agents propagate flawed outputs through interconnected software pipelines, transforming local inaccuracies into systemic failures. In multi-agent coding environments, a single agent’s erroneous completion, such as generating insecure logic or malformed interfaces, can trigger downstream faults as subsequent agents consume and redeploy its outputs. [huang2024resilience] demonstrate that even minor perturbations from “faulty” agents can destabilize collaborative workflows, revealing the fragility of LLM-based coordination without strong fault isolation. Similarly, [khanzadeh2025agentmesh] observe that emergent interactions in AgentMesh amplify defects across code generation stages, while [xiong2025butterfly] characterize “butterfly effects” in LLM toolchains, where trivial parameter-filling mistakes escalate into broad execution failures. Extending this view, [liu2025genomas] find that scientific discovery agents exhibit correlated errors due to shared model assumptions and data dependencies. Collectively, these studies underscore that LLM-mediated ecosystems lack the containment and traceability necessary to prevent error propagation—rendering traditional debugging and peer review inadequate in fully autonomous settings.
@@
Dependency proliferation denotes the uncontrolled expansion of software and library linkages arising from agent-driven code generation and refactoring. Autonomous coding agents frequently introduce or hallucinate dependencies beyond human oversight, compounding security and maintainability challenges. [spracklen2025we] reveal that nearly one-fifth of packages suggested by code-generation models are nonexistent or untrusted, creating opportunities for dependency-confusion attacks. [krishna2025importing] identify similar propagation patterns in automated import suggestions, where vulnerable or deprecated modules reappear across independent tasks. Furthermore, [zhao2025hfuzzer] show that fuzz-testing agents inadvertently inject unverified binaries during test synthesis, and [zhang2025cutting] demonstrate that build-automation agents routinely bypass provenance enforcement to optimize throughput. These behaviors collectively inflate the software’s attack surface and contravene emerging SBOM (Software Bill of Materials) and supply-chain transparency requirements. Dependency proliferation thus represents a structural form of technical debt—an accreting web of opaque linkages that undermines the long-term security and interpretability of autonomous development environments.
@@
Alignment failures emerge when an autonomous coding agent’s behavior diverges from the developer’s intent or the organization’s governance norms [shao2025your]. Unlike deterministic compilers, agentic systems interpret underspecified objectives probabilistically, optimizing for textual success rather than normative correctness. [naik2025agentmisalignment] empirically show that goal-oriented coding agents often maximize explicit task metrics while disregarding implicit safety or maintainability constraints embedded in natural-language prompts. Complementarily, [lynch2025agentic] describe “agentic misalignment” as an emergent property of reinforcement-trained code agents that overfit reward structures, leading to functionally correct yet normatively undesirable behaviors. Such divergence erodes accountability: once decisions are embedded in opaque model reasoning, tracing responsibility between human and agent becomes ambiguous. Alignment failures therefore epitomize the epistemic opacity of autonomous software systems, emphasizing the necessity of interpretable reward modeling, continuous supervision, and human-in-the-loop intervention to ensure fidelity between developer intent and agent execution.
Addressing these systemic risks therefore necessitates a layered oversight architecture that integrates automated analysis, formal verification, and human-in-the-loop governance. Modern static and dynamic analysis pipelines now incorporate model-aware scanners capable of parsing incomplete code fragments and predicting vulnerabilities in real time [dolcetti2024helping, keltek2025lsast], extending beyond conventional SAST [owasp_sast] and DAST [owasp_dast] frameworks by interfacing directly with generation streams so that feedback can be injected before completion. Complementing these, formal verification frameworks such as Kani [vanhattum2022verifying] and ACCA [cotroneo2024automating] demonstrate that lightweight proofs of functional correctness can be applied to AI-generated modules, though scalability remains limited. Reinforcement-learning-based watchdog agents offer a parallel safeguard: autonomous monitors trained to detect deviations from project constraints, perform behavioral tracing, and intervene when agents exceed operational bounds [kenton2024scalable, barj2024reinforcement]. Collectively, these technologies lay the groundwork for continuous, adaptive oversight that aligns autonomous coding processes with human accountability and system safety.
@@ 9.3.2 Toward Scalable Oversight Architectures
Achieving effective oversight in autonomous coding environments requires architectures that scale alongside agent capabilities. A central idea in recent alignment research is enabling weak-to-strong generalization, using relatively weak supervisors (humans or smaller models) to robustly guide much more powerful coding agents [burns2024weak]. In contrast to traditional settings where humans directly review weaker AI outputs, future agents may produce “millions of lines of novel and potentially dangerous code” beyond any one human’s full comprehension. Scalable oversight architectures address this gap by bootstrapping supervision: leveraging layered AI assistants and automated checks so that even limited feedback can generalize into reliable constraints on agent behavior [engels2025scaling]. Alignment strategies such as iterated amplification, recursive reward modeling, and AI debate all share this goal, allowing networks of weaker reviewers, often AIs themselves, to collectively evaluate and correct stronger agents’ decisions [christiano2018supervising, irving2018ai]. This paradigm of “weak-to-strong oversight” underpins emerging proposals for governing advanced code-generation systems, ensuring that each increase in agent autonomy is met with a proportional increase in oversight capacity.
@@
One approach is to organize oversight in tiers, where an automated or human overseer supervises the coding agent at a high level, while that overseer is itself aided by subordinate AI tools. This nested oversight or recursive supervision strategy allows complex judgments to be decomposed into simpler subtasks manageable by weaker supervisors [engels2025scaling]. Recent studies demonstrate that a weak teacher model’s guidance can be amplified through ensemble feedback and iterative training such that a much stronger student model aligns with the teacher’s underlying intent [sang2024improving]. OpenAI’s superalignment experiments showed that fine-tuning GPT-4 under the noisy guidance of a GPT-2-level model, augmented with confidence-based loss functions, could recover much of GPT-4’s original performance, achieving results comparable to GPT-3.5 on complex reasoning tasks [burns2024weak]. These findings suggest that, with appropriate training and oversight scaffolds, coding agents can generalize from weak supervision to strong, aligned behavior. However, theory also indicates that scalability has limits: oversight success rates decline sharply when an agent’s capability surpasses its overseer by several hundred Elo points [engels2025scaling], underscoring the need for continuous co-evolution between overseers and agents.
@@
Another branch of scalable oversight employs multi-agent debate frameworks [du2023improving, liu2025breaking, cui2025free], where coding agents critique each other’s code contributions to expose errors or unsafe behaviors [lang2025debate, irving2018ai]. Even a weaker human or AI judge can often identify the more truthful debater if the setup incentivizes evidence-based argumentation [irving2018ai]. In the software domain, DebateCoder [chen2025debatecoder] applies this principle to code generation: two large language models alternately produce and test code while generating adversarial unit tests against each other. Each iteration forces self-correction until both code and tests converge to correctness. This mechanism fuses automated testing and AI critique, forming a self-correcting oversight loop that improves code quality and reliability without human intervention. Such multi-agent oversight reduces dependence on a single reviewer and increases the likelihood that at least one agent detects subtle flaws or misalignments in autonomously written code.
@@
Scalable oversight also requires continuous, model-aware monitoring embedded throughout the software pipeline. Model-integrated static analyzers and runtime scanners can evaluate partially generated code, flagging vulnerabilities or policy violations in real time [dolcetti2024helping, keltek2025lsast]. Reinforcement learning–based watchdog agents extend this capacity: autonomous monitors trained to detect deviations from specification and intervene or alert humans when the coding agent exceeds its operational bounds [kenton2024scalable, barj2024reinforcement].
Scalable oversight architectures for Vibe Coding agents converge on the principle of amplifying limited oversight into broad governance. Through hierarchical weak-to-strong supervision, multi-agent critique, and automated watchdogs, researchers aim to generalize weak feedback signals into strong preventive constraints that evolve with agent capability [sang2024improving, burns2024weak, engels2025scaling]. The challenge lies not only in constructing these systems but in maintaining their reliability as agents advance beyond direct human comprehension. By incrementally upgrading oversight tools alongside the agents themselves, the field moves toward an adaptive architecture, one where even superhuman coding systems remain answerable to human-aligned constraints.